### v1

@@ 5 Conclusion
In the supplementary material, we first report additional experiments in Sec. F. We further show an additional visualization in Sec. G, and discuss the responsibility to human subjects and ethical consideration in Sec. H and I. Limitations and future work are summarized in Sec. J. A supplementary video is also provided as an additional illustration.
@@ F.2 Extension
In Table 7, the results show that our models all perform better than the baselines while using the same extractors. Notably, GeneFace has further used the HuBERT features to pre-train an intermediate representation on a large audio-visual corpus to enhance its generalizability for cross-domain audios. The generated results have been provided in the supplementary video.
@@
G Additional Visualization
@@ I Ethical Consideration
J Limitations and Future Work

### v2

@@ 5 Conclusion
In the supplementary material, we first report additional experiments in Sec. F and implementation details in Sec. G. We further show an additional visualization in Sec. H, and discuss the responsibility to human subjects and ethical consideration in Sec. I and J. Limitations and future work are summarized in Sec. K. A supplementary video is also provided as an additional illustration.
@@ F.2 Extension
In Figure 7, the results show that our models all perform better than the baselines while using the same extractors. Notably, GeneFace has further used the HuBERT features to pre-train an intermediate representation on a large audio-visual corpus to enhance its generalizability for cross-domain audios. The generated results have been provided in the supplementary video.
@@
G Implementation Details
G.1 Model Details.
Preprocessing. In the main paper, we use a frozen DeepSpeech [16] model to extract raw audio features. Then a CNN-based attention module in previous NeRF-based works [15, 40, 43, 24] is adopted to adapt and smooth the audio features. The upper-face expression feature is composed of a set of action units that only relate to the upper-face motions, specifically: 1, 2, 4, 5, 6, 7, and 45. We use the action units detected by OpenFace [3, 2] as the input signals. Utilizing the same preprocessor in previous works [43, 24], we first estimate the head pose via a BFM [35] facial model, and then inversely calculate the camera pose.
Persistent Gaussian Fields. The Persistent Gaussian Fields are built on the 3DGS repository  [^2] https://github.com/graphdeco-inria/gaussian-splatting. We init the fields with a random point cloud, and inherit the adaptive density control from 3DGS during training. For rendering, we adopt a modified 3DGS rasterizer  [^3] https://github.com/ashawkey/diff-gaussian-rasterization, which enables alpha supervision, to penalize the empty areas.
Grid-based Motion Fields. In the experiments, we use three 2D hash-encoders [32] with a 3-layer MLP decoder to implement the Grid-based Motion Fields. For each 2D hash-encoder, we set the resolution range from 16 to 256 in the face branch, and 64 to 384 for the inside mouth branch. All these encoders are set with 12 levels. The hidden dimension of the MLP decoder is set to 64 and 32, respectively for the face and inside mouth branches.
Optimizer. During training, we keep two optimizers to optimize the Persistent Gaussian Fields and Grid-based Motion Fields separately. For the Persistent Gaussian Fields, we inherit the Adam optimizer from 3DGS with a similar hyperparameter setting to optimize the Gaussians. For the motion fields, we adopt an AdamW optimizer with learning rates of $35e-$ 3 and $45e-$ 4 for the hash encoder and other parts. An exponential scheduler is used to adjust the learning rates.
G.2 Face-Mouth Segmentation
As mentioned in Sec. 3.3, a semantic mask is used to divide the face and inside mouth in the 2D image. Specifically, we use a combination of two off-the-shelf face parsing models to generate the mask. First, we use a BiSeNet parser [54] pre-trained on CelebAMask-HQ dataset [23] to predict a coarse mask that contains a coarse mouth segmentation for the whole head and inside mouth. Due to the domain gap, the mouth mask from this face parser may fail to completely cover the mouth sometimes. To enhance the robustness, we further introduce a ResNet-based FPN trained on EasyPortrait [19] as a tooth parser to predict a tooth mask. Finally, we overlay these two masks to get a finer one, then use the corresponding masked image as $\mathcal{I}_{\text{mask}}$ in Eq. (10, 11) to supervise the two branches of TalkingGaussian. For each branch, we apply the semantic mask on both the ground truth image and the predicted result during training. The overall pipeline is shown in Fig 8.
Figure 8: Illustration of face and inside mouth segmentation.
H Additional Visualization
@@ I Dataset Declaration
K Limitations and Future Work